# Towards White Box Deep Learning

### v2

@@ Abstract
This paper introduces semantic features as a candidate conceptual framework for building inherently interpretable neural networks. A proof of concept model for informative subproblem of MNIST consists of 4 such layers with the total of ~5K learnable parameters. The model is well-motivated, inherently interpretable, requires little hyperparameter tuning and achieves human-level adversarial test accuracy - with no form of adversarial training! These results and the general nature of the approach warrant further research on semantic features. The code is available at https://github.com/314-Foundation/white-box-nn.
@@ 1 Introduction
Machine learning systems used in safety-critical applications such as healthcare, military, autonomous vehicles or cybersecurity need to be reliable, i.e. they cannot fail unpredictably. The black box nature of state of the art (SOTA) deep neural networks (DNN) make it exceedingly hard to give such guarantees, which limits their usability to non-critical domains. The recent research in Explainable AI (XAI) focuses predominantly on post-hoc attribution methods which may provide explanations but, due to the complexity of SOTA, are hard to be converted to reliable fail-safes. This is a burning issue because of the wide susceptibility of DNNs to distribution shifts and adversarial attacks [18, 5] meaning that models can make arbitrarily incorrect predictions and their reported high test-time metrics don’t guarantee robust real-life performance111models often rely on spurious features that leak from test to train [6].. The recently popular Neurosymbolic AI [17] consists of higher-level methods that focus on augmenting neural network representations (machine perception) with background structured knowledge (reasoning) but don’t address the potential failures of the former.
@@
Existing methods of improving DNN robustness are far from ideal. Adversarial Training [11] being one of the strongest empirical defenses against attacks [1] is slow, data-hungry [15, 8] and results in robust generalisation gap [13, 21] as models overfit to the train-time adversaries. Thus, the biggest challenge of empirical approaches is reliable evaluation of robustness as more and more sophisticated attacks break the previously successful defenses [3, 4].
@@
To address these challenges the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [9]. However the exact certification is NP-Complete [20] and therefore has serious limitations in practice as it’s limited to relatively small problems and small networks and/or to probabilistic approaches. Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation. The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/. for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
The highlighted issues are relevant for datasets as simple as MNIST [16] or even some of its binary subsets [19] - despite being "toy" and almost linearly separable problems, they still remain elusive in terms of robustness, rising further questions about the true generalisation capabilities of SOTA as no neural network model can be honestly considered reliable and adequate333in terms of internal complexity; SOTA models define inappropriately complicated functions. for this allegedly "simple" problem.
1.1 Proposed approach
This paper proposes an intermediate approach to DNN reliability and introduces a candidate theoretical framework for building white box neural networks. Building and analysing a white box DNN would be similar to writing and debugging a procedural program - it doesn’t have to come with theoretical certificates but needs to be easily understandable and internalizable by the programmer. A good software has to be designed well enough to account for reasonable failure scenarios, handle foreseeable exceptions and, should an unforseen bug happen, be amenable for fixes. If the design and inner mechanics of a DNN are sufficiently clear then working with such model wouldn’t be very different from maintaining a standard application.
Current SOTA models are large, unwieldy black boxes that require prohibitive amounts of resources; yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135K neurons! This scale seems easily maintainable with proper abstractions. After all, the core algorithms used today are mostly unchanged since ~30 years! What if there really exists a better way?
@@ 1.1 Proposed approach
To find it we must start with the basics and take the agile approach. Focusing on the simplest relevant problem cuts off the obfuscating details and speeds up the experimentation phase. With that in mind the Minimum Viable Dataset (MVD) for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [19] and therefore might be considered harder than the entire MNIST in terms of average adversarial accuracy444the preliminary results on the full MNIST are discussed in Appendix C..
1.2 The core idea
This paper introduces semantic features as a general idea for sharing weights inside a neural network layer. The rigorous definition will be given in the next section. The concept is similar to that of "inverse rendering" in Capsule Networks [14] where features have many possible states and the best-matching state has to be found. Identifying different states by the subsequent layers gives rise to controlled dimensionality reduction. Thus semantic features aim to capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time. This is in fact a pretty strong regularization. As shown in this paper, choosing appropriate layers of semantic features for the MVD results in what can be considered as a white box neural network.
The remainder of the paper is structured as follows:
  • Section 2. gives a rigorous definition of semantic feature and argues for the generality of the notion;
  • Section 3. builds a carefully motivated PoC 4-layer white box neural network for selected MVD;
  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and ablation;
  • Section 5. provides a short discussion of the results;
The paper ends with ideas for further research and acknowledgments.
@@ 2 Semantic features
In machine learning inputs are represented as tensors. Yet the standard Euclidean topology in tensor space usually fails to account for small domain-specific variations of features. For example shifting an image by 1 pixel is semantically negligible but usually results in a distant output in $L_{2}$ metric.
@@
Intuitively a semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [14]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias555by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
@@ 2.1 Examples of semantic features
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel666can be made differentiable as in [12, 7].
@@
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices777preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images6
@@ 2.2 Matching semantic features
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates888i.e. $f_{p}$ is a conjunction of lower-level features.. Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks. This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
@@
  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 999alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
@@ 3 Network architecture
The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $5$K parameters.
@@ 3.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH101010the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@ 3.2 Convolutional Layer
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches111111a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
The layer is followed by ReLU activation. We initialize $g^{0}$ and $g^{1}$ as the identity kernel of shape $(1,1,5,5)$ plus Gaussian noise $\sim\mathcal{N}(0,0.1)$.
@@ 3.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small121212close to the identity augmented matrix affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@
For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is zero-padded evenly on every side and identified with the resulting $28\times 28$ image. Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
@@ 3.4 Logical Layer
After Affine Layer has extracted predictive shapes it remains to encode logical claims such as "number 3 is A or B and not C and not D" etc. Since the previous layer is learnable we can enforce a preferable structure on the input neurons to the Logical Layer. This layer will consist of 2 logical $f_{\mathbf{P}\mathbf{L}}$ (one for every label) defined as follows:
@@ 4.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability131313mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
@@
Figure 1: (left) Reliability test-time curve for the classic 40-step PGD Attack. (right) Learning curve.

### v4

@@ Abstract
This paper introduces semantic features as a candidate conceptual framework for white box neural networks. The proof of concept model is well-motivated, inherently interpretable, has low parameter-count and achieves almost human-level adversarial test metrics - with no adversarial training! These results and the general nature of the approach warrant further research on semantic features. The code is available at https://github.com/314-Foundation/white-box-nn.
@@ 1 Introduction
The core advantages of deep neural networks (DNNs) are their architectural simplicity and automatic feature learning. The latter is crucial for working with unstructured data as developers don’t need to design features by hand. However giving away the control over features leads to black box models - DNNs tend to learn uninterpretable "shortcut" correlations [13] that leak from train to test [15], hampering alignment and out-of-distribution performance. In particular, this gives rise to adversarial attacks [30] - semantically negligible perturbations of data that arbitrarily change model’s predictions. Adversarial vulnerability is a widespread phenomenon (vision, segmentation, NLP, sound, RL, tabular data) and largely contributes to the general lack of trust in DNNs, substantially limiting their adoption in high-stakes applications such as healthcare, military, autonomous vehicles or cybersecurity. Conversely, the main advantage of hand-designed features is the fine-grained control over model’s performance; however, such systems quickly become infeasibly complex.
This paper aims to address those issues by reconciling Deep Learning with feature engineering - with the help of locality engineering. Specifically, semantic features are introduced as a general conceptual machinery for controlled dimensionality reduction inside a neural network layer. Figure 1 presents the core idea behind the notion and the rigorous definition is given in Section 4. Implementing a semantic feature predominantly involves encoding appropriate invariants (i.e. semantic locality) for the desired semantic entities, e.g. "small affine perturbations" for shapes or "adding small $\delta\in[-\epsilon,\epsilon]$" for numbers. Thus, semantic features capture the core characteristic of real-world objects: having many possible states but being at exactly one state at a time. It turns out that for the trained proof-of-concept (PoC) DNN such parameter-sharing strategy leads to human-aligned base features - which motivates calling the PoC a white box neural network, see Section 6. In particular, the resulting inherent interpretability of the PoC allowed for designing a strong architectural defense111with no form of adversarial training! against adversarial attacks, as discussed in Section 7. Thus, the notion of "semantic feature" and the adversarially robust white box PoC network are the main contributions of this work.
Figure 1: Visualisation of semantic feature $f_{\mathbf{P}\mathbf{L}}$ and its matching mechanism (SFmatch, see Section 4.2). The semantic feature $f_{\mathbf{P}\mathbf{L}}$ consists of a base feature $f$ and a set of its "small" modifications $\mathbf{L}(p_{i})(f)$. In this case, the base feature has the same dimensions as the input image and its modifications are 2D affine transformations. $\mathrm{SFmatch}(d,f_{\mathbf{P}\mathbf{L}})$ takes the maximum of the dot products $<d,\mathbf{L}(p_{i})(f)>$ thus identifying all $\mathbf{L}(p_{i})(f)$ with $f$. Affine transformations are typically parameterized by extended matrices of dimension $2\times 3$; both the base feature and the parameters $p_{i}$ of its modifications are learned. If $f$ and its modifications are easily understood by humans, then the neural network layer composed of such features can be considered a white box layer. The precise definition of semantic feature is found in Section 4.
The paper is organised as follows. Section 2 investigates related work on explainability and adversarial robustness. Section 3 briefly describes the adopted methodology. Section 4 gives the rigorous definition of semantic feature and argues for the generality of the notion. Section 5 builds a carefully motivated PoC 4-layer white box neural network for the selected problem. Section 6 analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and performs an ablation study. Section 7 provides a short discussion of advantages and limitations of the approach. The paper ends with ideas for further research and acknowledgments.
2 Related work
2.1 Adversarial robustness
@@
Existing methods of improving DNN robustness are far from ideal. Adversarial Training [22] being one of the strongest empirical defenses against attacks [3] is slow, data-hungry [26, 18] and results in robust generalisation gap [24, 34] as models overfit to the train-time adversaries. This highlights the biggest challenge of empirical approaches, i.e. reliable evaluation of robustness, as more and more sophisticated attacks break the previously successful defenses [7, 9].
@@
To address this issue the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [19]. However the complete certification faces serious obstacles as it is NP-Complete [33] and therefore limited to small datasets and networks, and/or to probabilistic approaches. Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation. The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/. for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
2.2 Explainability
Recent work indicates a link between adversarial robustness and inherent interpretability as adversarially trained models tend to enjoy Perceptually-Aligned Gradients (PAG) - taking gradient steps along the input to maximise the activation of a selected neuron in the representation space often yields visually aligned features [10, 17]. There is also some evidence for the opposite implication that PAG implies certain form of adversarial robustness [12, 29]. However Section 6.4 of this paper shows that the latter implication does not hold - model can be inherently interpretable but not robust.
Susceptibility to adversarial attacks is arguably the main drawback of known approaches to Explainable AI (XAI). One would reasonably expect explainable models to rely exclusively on human-aligned features instead of spurious "shortcut" signals [13] and therefore be robust to adversaries. However the post-hoc attribution methods, which aim to explain the decisions of already-trained models, are fragile [1] and themselves prone to attacks [14]. On the other hand the inherently interpretable solutions such as Prototypical Parts [20, 8], Bag-of-Features [5], Self-Explaining Networks [2], SITE [32], Recurrent Attention [11], Neurosymbolic AI [28] or Capsule Networks [25] all employ various classic black box encoder networks and thus introduce the adversarial vulnerability into their cores.
Overall the field of adversarially robust explanations (AdvXAI) is fairly unexplored; for more information on this emerging research domain see the recent survey [4].
3 Methodology
@@
This paper focuses on the simplest relevant problem (Minimum Viable Dataset, MVD) to cut off the obfuscating details and get into the core of what makes neural networks fragile and uninterpretable. Despite much effort, DNNs remain susceptible to adversarial attacks even on MNIST [27] and therefore cannot be considered human-aligned, see Figure 2. That being said, the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [31] and might be considered harder than the entire MNIST in terms of average adversarial accuracy. For completeness, the preliminary results on the full MNIST are discussed in Appendix C.
The paper provides both qualitative and quantitative results. The qualitative analysis focuses on visualising the internal distribution of weights inside layers and how features are matched against the input. For the quantitative results a strong adversarial regime is employed besides computing the standard clean test metrics (computing the clean metrics alone is considered a flawed approach for the reasons discussed earlier).
Figure 2: A typical DNN for MNIST can be arbitrarily fooled by adding semantically negligible noise.
@@ 4 Semantic features
In machine learning inputs are represented as tensors. Yet, the standard Euclidean topology in tensor space usually fails to account for small domain-specific variations of features. For example shifting an image by 1 pixel is semantically negligible but usually results in a distant output in $L_{2}$ metric.
@@
Intuitively, semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [25]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias333by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
@@ 4.1 Examples of semantic features
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel444can be made differentiable as in [23, 16].
@@
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices555preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images4
@@ 4.2 Matching semantic features
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates666i.e. $f_{p}$ is a conjunction of lower-level features.. Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks. This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
@@
  • for convolutional777an alternative definition would treat every $g^{i}$ as a separate semantic feature; we stick to the max option for simplicity. $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$
@@ 5 Network architecture
The network consists of the following 4 layers stacked sequentially. For simplicity we don’t add bias to any of those layers. Te resulting model has around $5$K parameters. The architecture is visualized in the Figure 3.
Figure 3: Architecture of the PoC white box neural network. The first two layers consist of semantic features that operate per-pixel - the first layer takes into account only the pixel’s value, while the second examines its $5\times 5$ neighborhood to determine if the pixel lays on a "bright line". The first two layers retain the shape of the input image. The third layer comprises of 8 affine $f_{\mathbf{P}\mathbf{L}}$ as visualized in Figure 1. The final layer consists of two logical $f_{\mathbf{P}\mathbf{L}}$: intuitively, the first one checks whether at least one affine $f_{\mathbf{P}\mathbf{L}}$ corresponding to "3" is active and none of the affine $f_{\mathbf{P}\mathbf{L}}$ corresponding to "5" are active; the second logical $f_{\mathbf{P}\mathbf{L}}$ works in the opposite way. Section 5 describes the architecture in more detail.
@@ 5.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH888the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@ 5.2 Convolutional Semantic Layer
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with the $5\times 5$ neighbourhood of $(x,y)$ and takes the maximum across all the matches999a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively, this layer checks if pixels have adequately structured neighbourhoods.
The layer is followed by ReLU activation. We initialize $g^{0}$ and $g^{1}$ each as the identity kernel of shape $(1,1,5,5)$ plus Gaussian noise $\sim\mathcal{N}(0,0.1)$.
@@ 5.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small101010close to the identity augmented matrix. affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@
For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is zero-padded evenly on every side and identified with the resulting $28\times 28$ image. Intuitively, this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
@@ 5.4 Logical Layer
After Affine Layer has extracted predictive shapes it remains to encode logical claims such as "number 3 is A or B and not C and not D" etc. Since the previous layer is learnable, we can enforce a preferable structure on the input neurons to the Logical Layer. This layer will consist of 2 logical $f_{\mathbf{P}\mathbf{L}}$ (one for every label) defined as follows:
@@ 6.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability111111mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
@@
Figure 4: (left) Reliability test-time curve (see Section 6.2 for details). (right) Learning curve. Both test and validation metrics presented here are computed under the classic 40-step PGD Attack.